At what annual rate compounded continously will $1000 have to be invested to amount to $2,500 in 10 years?
Answer by stanbon(75887)

At what annual rate compounded continously will $1000 have to be invested to amount to $2,500 in 10 years?
------
A(t) = Ao*e^(rt)
----
2500 = 1000*e^(10r)
e^(10r) = 2.5
----
10r = ln(2.5)
r = (1/10)*ln(2.5)
---
r = 0.0916 ~ 9%
===================
cheers,
Stan H.
===================